For electromagnetic radiation with a wavelength of 213.5 nm: (a) What is the frequency of the radiation (in s-1)? (b) What is the energy (in J) of one photon of the radiation? (c) What is the energy (in kJ) of one mole of photons of the radiation?
Answer #1

E= h\nu

Where h is planks constant.

and \nu is frequency.

1. a) Given Wavelength = 213-5 nm As c= 8X, where c= Velocity of light = frequency x= Wareleufth. 3X10% m/s - V x 213-5x169m5 As I mol of photons = 6.023x102% pho ton As Energy of I photon is 9.310 x 10 195 Hence Energy of 1 meal of photon which is